Calculate the frequency, phase velocity and maximum particle velocity for y=0.2sin (0.8x-40t)
1
Expert's answer
2021-09-24T09:25:56-0400

(a)

"\omega=2\pi f,""f=\dfrac{\omega}{2\pi}=\dfrac{40\ \dfrac{rad}{s}}{2\pi}=6.37\ Hz."

(b)

"v_{p} = \dfrac{\omega}{k}=\dfrac{40\ \dfrac{rad}{s}}{0.8\ m^{-1}}=50\ \dfrac{m}{s}."

(c) The particle velocity can be found as follows:


"v=\dfrac{dy}{dt}=-8cos(0.8x-40t)."

The maximum particle velocity will be when "cos(0.8x-40t)=\pm1". Therefore, the magnitude of the maximum particle velocity equals "v_{max}=8\ m/s."

Answer:

(a) "f=6.37\ Hz."

(b) "v_{p} =50\ \dfrac{m}{s}."

(c) "v_{max}=8\ m/s."